# ALMA FACTS III. High-Resolution CO(2-1)/CO(1-0) Maps of Twelve Nearby Galaxies

### v1

@@ 4.1.5 NGC 3521
fluctuates between low and high values along the spiral arms (Figure 7). Low-ratio and high-ratio encompass 53% and 45% of the disk respectively (blue histogram). The CO(1-0) intensity contained by the low-ratio and high-ratio gas are 50% and 49% respectively (green histogram). Over the whole area, the mean is 0.69 while the intensity-weighted mean is 0.70. As both transitions of CO are not detected in the center, there is little contribution from the central kiloparsec ( kpc) in the distributions, if any.
4.1.6 NGC 3621
@@
From Figure 3, the molecular gas structure appears less organized and flocculent. fluctuates between low and high values across the disk (Figure 7). The spatial distribution of high-ratio gas seems to coincide closely with HII regions. Low-ratio gas and high-ratio gas cover 36% and 57% of the area shown (blue histogram). The CO(1-0) intensity contained by low-ratio and high-ratio gas are 27% and 67% (green histogram). Hence, the majority of the molecular gas by area and by CO(1-0) intensity have high-ratios. Over the whole area, the mean is 0.75 while the intensity-weighted mean is 0.78.
4.1.7 NGC 3627
@@
is high in the center, lowers along the bar, and then increases again at the bar ends (Figure 8). The high ratio regions especially at the bar ends and along the spiral arms are located near the HII regions in the optical image. Low-ratio and high-ratio gas encompass 59% and 36% of the galaxy by area (blue histogram), while they contain a similar amount of the CO(1-0) intensity at 49% and 49% (green histogram), respectively. Over the whole area, the mean is 0.68 while the intensity-weighted mean is 0.71. The majority of the gas in the central kiloparsec ( kpc) is high-ratio gas.
den Brok et al. (2023) also studied in NGC 3627 (as well as other line ratios including CO(3-2) ALMA observations) at resolution (200 pc). They also found high in the center and bar-ends, in comparison to the bar and spiral arms.
4.1.8 NGC 4254
NGC 4254 is an interacting spiral galaxy (SAc) in the Virgo cluster. The angular resolution of 2.5 is 168 pc. From Figure 4, the molecular gas distribution appears asymmetric (e.g., Hidaka & Sofue, 2002). The center is bright and spiral arms are apparent in both transitions of CO, but the structures of the spiral arm and tightness differ between the northern and southern parts. The southern spiral arm extends beyond the coverage in CO (e.g., Figure 1).
fluctuates between low and high-values along the spiral arms in the disk. Low-ratio and high-ratio gas encompass 52% and 44% of the area (blue histogram). Low-ratio and high-ratio gas include similar portions of the CO(1-0) intensity distribution, at 46% and 52% respectively (green histogram). Over the whole area, the mean is 0.69 while the intensity-weighted mean is 0.71. From the histogram, the central kiloparsec ( kpc) contains mostly high-ratio gas.
4.1.9 NGC 4321
NGC 4321 is a barred spiral galaxy (SABbc). The 2.5 resolution corresponds to 173 pc. From Figure 4, the molecular gas distribution shows a bright center and two spiral arms in both CO transitions. The spiral arms extend beyond the CO coverage (e.g., Figure 1).
@@ 4.1.9 NGC 4321
is high in the center, lowers in the bar region, but fluctuates between low and high-values along the spiral arms (Figure 9). The southern spiral arm may be showing a potential transition from low to high-ratio gas along the gas flow across the spiral arm (where it runs horizontally in the image), assuming that the gas passes through the arm in the counter-clockwise direction. Low-ratio gas encompasses 63% of the disk while high-ratio gas makes up 30% by area (blue histogram). Likewise, low-ratio gas contains 55% of the CO(1-0) intensity and high-ratio gas contains 39% (green histogram). Most of the molecular gas distribution in NGC 4321 thus has low-ratios by area and by CO(1-0) intensity. Over the whole area, the mean is 0.64 while the intensity-weighted mean is 0.67. The central 1-2 kpc is mostly high ratio gas.
4.1.10 NGC 4536
@@
is high in the center but fluctuates between low and high-values along the spiral arms (Figure 9). The peak of the distribution by number is in low-ratio gas (blue histogram), but the peak shifts to high-ratio gas when weighted by CO(1-0) intensity (green histogram). Low-ratio gas and high-ratio gas encompass 70% and 27% of the area. However, low-ratio gas contains 32% of the CO(1-0) intensity distribution, with high-ratio gas containing 52%. Over the whole area, the mean is 0.61 while the intensity-weighted mean is 0.79. From the histograms, there is molecular gas concentrated in the central region, with the central 1-2 kpc containing predominantly high/very-high ratio gas.
4.1.11 NGC 4579
NGC 4579 is a barred spiral galaxy (SABb). At its distance, the angular resolution of 2.5 corresponds to a spatial resolution of 199 pc. It is the farthest galaxy in the sample. From Figure 4 the central region is bright in both CO transitions and is surrounded by spiral arms. There is a concentration of gas in the central region. Its structure is spatially resolved and in the CO(1-0) map, it also looks like the roots of offset ridges. We treat it as a central concentration but since the transition from a central concentration to offset ridges is continuous, the boundary cannot be defined clearly.
is high in the center. The central gas concentration and high ratio region is also larger than the nominal definition of the central 1 kpc. Indeed, the histogram for the central 2 kpc explains most of the HRG. appears to fluctuate along the spiral structure (Figure 10). The distribution of the in the histogram weighted by CO(1-0) intensity appears doubly-peaked, signifying the difference in the distributions between the central region and spiral arms structure. The peaks of the distributions by number and when weighted by CO(1-0) are both in low-ratio gas. Low-ratio gas and high-ratio gas encompass 88% and 10% of the area, respectively (blue histogram). However low-ratio gas contains 74% of the CO(1-0) intensity while high-ratio gas contains 16%. Over the whole area, the mean is 0.53 while the intensity-weighted mean is 0.62. From the histograms, the central 1-2 kpc contains mostly high/very-high ratio gas, while the spiral arms contain low-ratio gas.
4.1.12 NGC 4826
@@
is high in the center but fluctuates between low and high in the surrounding regions. Low-ratio and high-ratio gas comprise similar portions of the galaxy by area, at 48% and 49% respectively. However, low-ratio gas contains 36% of the CO(1-0) intensity, while high-ratio gas contains 56%. Over the whole area, the mean is 0.71 while the intensity-weighted mean is 0.76. Given the small linear extent of the molecular gas, this skewed distribution towards high-ratio gas may be consistent with those in the central regions of the other galaxies, especially in the barred galaxies (see Figure 13). This might indicate that once the gas is concentrated in the central regions, becomes high independent of the cause of the concentration.
(1) | (2) | (3) | (4) | (5) | |
Galaxy | FoV | Beam Size | RMS Noise (1) | ||
Map Area, P.A. | bmaj, bmin, P.A. | Native Res. | Smoothed (2.5) | ||
[, ] | [, , ] | [mK] | [mK] | ||
CO(2-1) Data Cube Parameters | |||||
NGC 0628 | 4.3 2.9, 135 | 1.48, 0.88, -61.04 | 74.9 | 24.9 | |
NGC 1097 | 4.4 2.5, 145 | 1.54, 1.17, -85.48 | 46.2 | 18.5 | |
NGC 1512 | 3.1 2.2, 37 | 1.23, 0.99, 85.14 | 71.9 | 21.8 | |
NGC 3351 | 2.6 2.5, 0 | 1.44, 1.20, -78.83 | 92.2 | 36.2 | |
NGC 3521 | 4.9 2.1, 160 | 1.39, 1.12, -81.01 | 37.9 | 13.6 | |
NGC 3621 | 4.2 2.3, 160 | 1.55, 1.13, -79.09 | 31.8 | 12.2 | |
NGC 3627 | 4.2 2.3, 0 | 1.16, 1.00, 44.23 | 92.0 | 28.1 | |
NGC 4254 | 3.5 3.0, 0 | 1.64, 1.17, 57.54 | 53.1 | 22.4 | |
NGC 4321 | 3.6 3.1, 90 | 1.13, 1.00, -30.25 | 102.5 | 28.5 | |
NGC 4536 | 4.2 1.8, 119 | 1.55, 1.22, -71.86 | 25.3 | 10.0 | |
NGC 4579 | 3.0 2.2, 52 | 1.52, 1.33, -73.31 | 39.3 | 16.8 | |
NGC 4826 | 2.8 1.8, 112 | 1.35, 1.18, -23.85 | 53.6 | 19.1 | |
CO(1-0) Data Cube Parameters | |||||
NGC 0628 | 4.3 3.3, 135 | 2.42, 1.94, -56.77 | 63.8 | 49.0 | |
NGC 1097 | 4.8 2.8, 145 | 2.11, 1.69, 89.44 | 75.7 | 47.4 | |
NGC 1512 | 3.1 2.5, 37 | 2.24, 1.70, 88.36 | 60.6 | 39.9 | |
NGC 3351 | 2.9 2.8, 0 | 2.10, 2.02, -72.57 | 60.0 | 42.0 | |
NGC 3521 | 5.3 2.5, 160 | 2.18, 1.95, 88.84 | 67.1 | 47.5 | |
NGC 3621 | 4.4 2.9, 160 | 2.16, 1.64, 84.68 | 75.5 | 47.0 | |
NGC 3627 | 4.4 2.9, 0 | 2.17, 2.08, -46.96 | 68.8 | 50.4 | |
NGC 4254 | 4.0 3.7, 0 | 2.41, 1.95, -49.85 | 68.9 | 52.8 | |
NGC 4321 | 4.0 3.6, 90 | 2.17, 1.95, -58.89 | 68.7 | 48.2 | |
NGC 4536 | 4.5 2.2, 119 | 2.10, 2.01, 85.63 | 59.2 | 41.3 | |
NGC 4579 | 3.1 2.6, 52 | 2.20, 1.93, -51.86 | 74.0 | 52.2 | |
NGC 4826 | 3.1 2.1, 112 | 2.45, 2.02, -22.37 | 55.1 | 44.4 |
@@
Note. — (1) Galaxy (2) Rectangular field-of-view (FoV) enclosing the region out to where the sensitivity drops to 50% of the peak sensitivity (40% for NGC 4321 CO(2-1)). The position angle (PA) is along the major axis of the map, counterclockwise from North. (3) Beam Size. (4)(5) RMS Noise for the 12m+07m+TP data in 5 km/s channel at the native and smoothed resolutions, respectively.
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | |
Galaxy | by Area | by | |||||||
[M⊙ pc-2] | [M⊙ pc-2] | Mean | LRG [%] | HRG [%] | Mean | LRG [%] | HRG [%] | ||
NGC0628 | 40 | 34 | 0.62 | 70 | 24 | 0.59 | 78 | 18 | |
NGC1097 | 647 | 71 | 0.61 | 73 | 24 | 0.70 | 48 | 50 | |
NGC1512 | 61 | 12 | 0.67 | 58 | 36 | 0.76 | 37 | 56 | |
NGC3351 | 319 | 18 | 0.65 | 67 | 30 | 0.73 | 44 | 55 | |
NGC3521 | 24 | 40 | 0.69 | 53 | 45 | 0.70 | 50 | 49 | |
NGC3621 | 21 | 16 | 0.75 | 36 | 57 | 0.78 | 27 | 67 | |
NGC3627 | 383 | 39 | 0.68 | 59 | 36 | 0.71 | 49 | 49 | |
NGC4254 | 266 | 65 | 0.69 | 52 | 44 | 0.71 | 46 | 52 | |
NGC4321 | 490 | 67 | 0.64 | 63 | 30 | 0.67 | 55 | 39 | |
NGC4536 | 403 | 19 | 0.61 | 70 | 27 | 0.79 | 32 | 52 | |
NGC4579 | 240 | 30 | 0.53 | 88 | 10 | 0.62 | 74 | 16 | |
NGC4826 | 200 | 61 | 0.71 | 48 | 49 | 0.76 | 36 | 56 |
@@
(4-6) Mean , and percentage of molecular gas in LRG and HRG across the whole area of the galaxy.
(7-9) Same as (4-6) but weighted by .
5 Discussion

### v2

@@ IV.1.5 NGC 3521
fluctuates between low and high values along the spiral arms (Figure 7). Low-ratio and high-ratio encompass 52% and 45% of the disk respectively (blue histogram). The CO(1-0) intensity contained by the low-ratio and high-ratio gas are 50% and 49% respectively (green histogram). Over the whole area, the mean is 0.69 while the intensity-weighted mean is 0.70. As both transitions of CO are not detected in the center, there is little contribution from the central kiloparsec ( kpc) in the distributions, if any.
IV.1.6 NGC 3621
@@
From Figure 3, the molecular gas structure appears less organized and flocculent. fluctuates between low and high values across the disk (Figure 7). The spatial distribution of high-ratio gas seems to coincide closely with HII regions. Low-ratio gas and high-ratio gas cover 35% and 58% of the area shown (blue histogram). The CO(1-0) intensity contained by low-ratio and high-ratio gas are 26% and 67% (green histogram). Hence, the majority of the molecular gas by area and by CO(1-0) intensity have high-ratios. Over the whole area, the mean is 0.76 while the intensity-weighted mean is 0.78.
IV.1.7 NGC 3627
@@
is high in the center, lowers along the bar, and then increases again at the bar ends (Figure 8). The high ratio regions especially at the bar ends and along the spiral arms are located near the HII regions in the optical image. Low-ratio and high-ratio gas encompass 57% and 38% of the galaxy by area (blue histogram), while they contain a similar amount of the CO(1-0) intensity at 48% and 50% (green histogram), respectively. Over the whole area, the mean is 0.70 while the intensity-weighted mean is 0.71. The majority of the gas in the central kiloparsec ( kpc) is high-ratio gas.
den Brok et al. (2023) also studied in NGC 3627 (as well as other line ratios including CO(3-2) ALMA observations) at resolution (200 pc), and correlated it with star-formation activity (H). They also found high in the center and bar-ends, in comparison to the bar and spiral arms.
IV.1.8 NGC 4254
NGC 4254 is an interacting unbarred spiral galaxy (SAc) in the Virgo cluster. The angular resolution of 2.5 is 168 pc. From Figure 4, the molecular gas distribution appears asymmetric (e.g., Hidaka & Sofue, 2002). The center is bright and spiral arms are apparent in both transitions of CO, but the structures of the spiral arm and tightness differ between the northern and southern parts. The southern spiral arm extends beyond the coverage in CO (e.g., Figure 1).
fluctuates between low and high-values along the spiral arms in the disk. Low-ratio and high-ratio gas encompass 46% and 49% of the area (blue histogram). Low-ratio and high-ratio gas include 44% and 54% of the CO(1-0) intensity distribution (green histogram), respectively. Over the whole area, the mean is 0.72 while the intensity-weighted mean is 0.71. From the histogram, the central kiloparsec ( kpc) contains mostly high-ratio gas.
IV.1.9 NGC 4321
NGC 4321 is a barred spiral galaxy (SABbc) in the Virgo cluster. The 2.5 resolution corresponds to 173 pc. From Figure 4, the molecular gas distribution shows a bright center and two spiral arms in both CO transitions. The spiral arms extend beyond the CO coverage (e.g., Figure 1).
@@ IV.1.9 NGC 4321
is high in the center, lowers in the bar region, but fluctuates between low and high-values along the spiral arms (Figure 9). The southern spiral arm may be showing a potential transition from low to high-ratio gas along the gas flow across the spiral arm (where it runs horizontally in the image), assuming that the gas passes through the arm in the counter-clockwise direction. Low-ratio gas encompasses 56% of the disk while high-ratio gas makes up 35% by area (blue histogram). Likewise, low-ratio gas contains 49% of the CO(1-0) intensity and high-ratio gas contains 44% (green histogram). Most of the molecular gas distribution in NGC 4321 thus has low-ratios by area and by CO(1-0) intensity. Over the whole area, the mean is 0.70 while the intensity-weighted mean is 0.71. The central 1-2 kpc is mostly high ratio gas.
IV.1.10 NGC 4536
@@
is high in the center but fluctuates between low and high-values along the spiral arms (Figure 9). The peak of the distribution by number is in low-ratio gas (blue histogram), but the peak shifts to high-ratio gas when weighted by CO(1-0) intensity (green histogram). Low-ratio gas and high-ratio gas encompass 68% and 28% of the area. However, low-ratio gas contains 31% of the CO(1-0) intensity distribution, with high-ratio gas containing 53%. Over the whole area, the mean is 0.62 while the intensity-weighted mean is 0.79. From the histograms, there is molecular gas concentrated in the central region, with the central 1-2 kpc containing predominantly high/very-high ratio gas.
IV.1.11 NGC 4579
NGC 4579 is a barred spiral galaxy (SABb). At its distance, the angular resolution of 2.5 corresponds to a spatial resolution of 199 pc. It is the farthest galaxy in the sample. From Figure 4, the central region is bright in both CO transitions and is surrounded by spiral arms. There is a concentration of gas in the central region. Its structure is spatially resolved and in the CO(1-0) map, it also looks like the roots of offset ridges. We treat it as a central concentration but since the transition from a central concentration to offset ridges is continuous, the boundary cannot be defined clearly.
is high in the center. The central gas concentration and high ratio region is also larger than the nominal definition of the central 1 kpc. Indeed, the histogram for the central 2 kpc explains most of the HRG. appears to fluctuate along the spiral structure (Figure 10). The distribution of the in the histogram weighted by CO(1-0) intensity appears doubly-peaked, signifying the difference in the distributions between the central region and spiral arms structure. The peaks of the distributions by number and when weighted by CO(1-0) are both in low-ratio gas. Low-ratio gas and high-ratio gas encompass 88% and 10% of the area, respectively (blue histogram). However low-ratio gas contains 74% of the CO(1-0) intensity while high-ratio gas contains 17%. Over the whole area, the mean is 0.54 while the intensity-weighted mean is 0.63. From the histograms, the central 1-2 kpc contains mostly high/very-high ratio gas, while the spiral arms contain low-ratio gas.
IV.1.12 NGC 4826
@@
is high in the center but fluctuates between low and high in the surrounding regions. Low-ratio and high-ratio gas comprise similar portions of the galaxy by area, at 47% and 49% respectively. However, low-ratio gas contains 36% of the CO(1-0) intensity, while high-ratio gas contains 56%. Over the whole area, the mean is 0.72 while the intensity-weighted mean is 0.76. Given the small linear extent of the molecular gas, this skewed distribution towards high-ratio gas may be consistent with those in the central regions of the other galaxies, especially in the barred galaxies (see Figure 13). This might indicate that once the gas is concentrated in the central regions, becomes high independent of the cause of the concentration.
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Galaxy | FoV | Beam Size | RMS Noise (1) | |||
| Map Area, P.A. | bmaj, bmin, P.A. | Native Res. | Smoothed (2.5) | |||
| [, ] | [, , ] | [mK] | [mK] | [M⊙ pc-2] | ||
| CO(2-1) Data Cube Parameters | ||||||
| NGC 0628 | 4.3 2.9, 135 | 1.48, 0.88, -61.04 | 74.9 | 24.9 | 0.8, 1.8 | |
| NGC 1097 | 4.4 2.5, 145 | 1.54, 1.17, -85.48 | 46.2 | 18.5 | 0.6, 1.3 | |
| NGC 1512 | 3.1 2.2, 37 | 1.23, 0.99, 85.14 | 71.9 | 21.8 | 0.7, 1.6 | |
| NGC 3351 | 2.6 2.5, 0 | 1.44, 1.20, -78.83 | 92.2 | 36.2 | 1.1, 2.6 | |
| NGC 3521 | 4.9 2.1, 160 | 1.39, 1.12, -81.01 | 37.9 | 13.6 | 0.4, 1.0 | |
| NGC 3621 | 4.2 2.3, 160 | 1.55, 1.13, -79.09 | 31.8 | 12.2 | 0.4, 0.9 | |
| NGC 3627 | 4.2 2.3, 0 | 1.16, 1.00, 44.23 | 92.0 | 28.1 | 0.9, 2.0 | |
| NGC 4254 | 3.5 3.0, 0 | 1.64, 1.17, 57.54 | 53.1 | 22.4 | 0.7, 1.6 | |
| NGC 4321 | 3.6 3.1, 90 | 1.13, 1.00, -30.25 | 102.5 | 28.5 | 0.9, 2.1 | |
| NGC 4536 | 4.2 1.8, 119 | 1.55, 1.22, -71.86 | 25.3 | 10.0 | 0.3, 0.7 | |
| NGC 4579 | 3.0 2.2, 52 | 1.52, 1.33, -73.31 | 39.3 | 16.8 | 0.5, 1.2 | |
| NGC 4826 | 2.8 1.8, 112 | 1.35, 1.18, -23.85 | 53.6 | 19.1 | 0.6, 1.4 | |
| CO(1-0) Data Cube Parameters | ||||||
| NGC 0628 | 4.3 3.3, 135 | 2.42, 1.94, -56.77 | 63.8 | 49.0 | 1.1 | |
| NGC 1097 | 4.8 2.8, 145 | 2.11, 1.69, 89.44 | 75.7 | 47.4 | 1.0 | |
| NGC 1512 | 3.1 2.5, 37 | 2.24, 1.70, 88.36 | 60.6 | 39.9 | 0.9 | |
| NGC 3351 | 2.9 2.8, 0 | 2.10, 2.02, -72.57 | 60.0 | 42.0 | 0.9 | |
| NGC 3521 | 5.3 2.5, 160 | 2.18, 1.95, 88.84 | 67.1 | 47.5 | 1.0 | |
| NGC 3621 | 4.4 2.9, 160 | 2.16, 1.64, 84.68 | 75.5 | 47.0 | 1.0 | |
| NGC 3627 | 4.4 2.9, 0 | 2.17, 2.08, -46.96 | 68.8 | 50.4 | 1.1 | |
| NGC 4254 | 4.0 3.7, 0 | 2.41, 1.95, -49.85 | 68.9 | 52.8 | 1.1 | |
| NGC 4321 | 4.0 3.6, 90 | 2.17, 1.95, -58.89 | 68.7 | 48.2 | 1.0 | |
| NGC 4536 | 4.5 2.2, 119 | 2.10, 2.01, 85.63 | 59.2 | 41.3 | 0.9 | |
| NGC 4579 | 3.1 2.6, 52 | 2.20, 1.93, -51.86 | 74.0 | 52.2 | 1.1 | |
| NGC 4826 | 3.1 2.1, 112 | 2.45, 2.02, -22.37 | 55.1 | 44.4 | 1.0 | |
@@
Note. — (1) Galaxy (2) Rectangular field-of-view (FoV) enclosing the region out to where the sensitivity drops to 50% of the peak sensitivity (40% for NGC 4321 CO(2-1)). The position angle (PA) is along the major axis of the map, counterclockwise from North. (3) Beam Size. (4) RMS Noise for the 12m+07m+TP data in 5 km/s channel at the native resolution. (5)(6) Same as (4) but at the smoothed 2.5 resolution and converted to mass surface density assuming = 4.35 and fiducial =0.7 and =0.3, respectively. The sensitivity limit becomes more important for lower brightness regions, which tend to have lower ratio.
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | |
| Galaxy | Ridge | by Area | by | |||||||
| [M⊙ pc-2] | [M⊙ pc-2] | Mean | LRG [%] | HRG [%] | Mean | LRG [%] | HRG [%] | |||
| NGC0628 | – | 40 | 36 | 0.65 | 67 | 26 | 0.60 | 77 | 19 | |
| NGC1097 | L | 647 | 75 | 0.63 | 70 | 26 | 0.71 | 47 | 50 | |
| NGC1512 | S | 61 | 13 | 0.72 | 50 | 42 | 0.79 | 31 | 61 | |
| NGC3351 | S | 319 | 22 | 0.69 | 58 | 38 | 0.77 | 36 | 62 | |
| NGC3521 | – | 24 | 41 | 0.69 | 52 | 45 | 0.70 | 50 | 49 | |
| NGC3621 | – | 21 | 17 | 0.76 | 35 | 58 | 0.78 | 26 | 67 | |
| NGC3627 | L | 394 | 41 | 0.70 | 57 | 38 | 0.71 | 48 | 50 | |
| NGC4254 | – | 266 | 70 | 0.72 | 46 | 49 | 0.71 | 44 | 54 | |
| NGC4321 | L | 490 | 77 | 0.70 | 56 | 35 | 0.71 | 49 | 44 | |
| NGC4536 | S | 403 | 20 | 0.62 | 68 | 28 | 0.79 | 31 | 53 | |
| NGC4579 | S | 240 | 31 | 0.54 | 88 | 10 | 0.63 | 74 | 17 | |
| NGC4826 | – | 200 | 62 | 0.72 | 47 | 49 | 0.76 | 36 | 56 | |
@@
(2) Long or short (“L” or “S”) offset ridges for barred galaxies, based on the molecular gas morphology (Section V.1, Figure 11).“-” stands for unbarred galaxies.
@@
(5-7) Mean , and percentage of molecular gas in LRG and HRG across the whole area of the galaxy.
(8-10) Same as (5-7) but weighted by .
V Discussion